Problem 4: Evaluation of the convolution integral too y(t) = (f * h)(t) = f(t)h(t – 7)dt is greatly simplified when either thh(t) - N h(t) . If (b) t f(t) h(t) (3)(3) (3) (C) 1 1 -2 -1 0 1 2Problem 5: This problem extends the method developed in the previous problem to cases where the derivative of either f(t) ord) Using the results in parts b and c, calculate y(t). e) It is not practical to experimentally determine the impulse respons
Answer #1

og 4@. fck) *564-1) = ? : 864-7) *hce) a hét-T) ? le have: 2,64) * X,Fe) = fx,(mazlt-podr * {(t-1)xh(t): face-T) hc4-mdr + f(@ fcel hrt) isi f)= -2864-1) ht) = $(t-1) + 8(t) +8C++) ->yet) = -2804-1) ¥h(t). ... | nyet) = f(4)*suck) -2h(t-) , = f(e) x[